OpenStudy (anonymous):

Find the volume of a regular tetrahedron whose altitude measures 10 cm.

hartnn (hartnn):

the formula is V = \(\sqrt 2\)/12 * a^3

OpenStudy (anonymous):

it is not the edge that is given but the ALTITUDE..

OpenStudy (anonymous):

now i got it.. :))

OpenStudy (anonymous):

How ?

OpenStudy (anonymous):

altitude = root(2/3) * a 10 = root(2/3) * a 100=2/3 a^2 a= 5 root 6 V = root(2)/12 * a^3 V = 216.5

OpenStudy (anonymous):

I thought it like this..\[\large\color{green}{x=\frac{100}{\sqrt{3}}}\] 

OpenStudy (anonymous):

Sorry..\[\large\color{green}{x=\frac{10}{\sqrt{3}}}\]

OpenStudy (anonymous):

that's fine.. ^^

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

@eroshea I am glad tht u got it..:D :)

OpenStudy (anonymous):

thanks for your help @hartnn and @Miyuru ^^

OpenStudy (anonymous):

I didn't help.. o.O

OpenStudy (anonymous):

Thank @hartnn ONLY

OpenStudy (anonymous):

well, you tried helping me.. so thanks :))

hartnn (hartnn):

u got it on your own, @eroshea :)

OpenStudy (anonymous):

Is my way wrong @hartnn

hartnn (hartnn):

if u got the same volume, then u are correct.

OpenStudy (anonymous):

It's nearly 181.4 :(

hartnn (hartnn):

then incorrect :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

OHH Anyway thnks for U all and Welcome...

OpenStudy (anonymous):

Ohh @eroshea you mean the height is 10.. Isn't it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I thought it was the altitude of a triangle.. :P

OpenStudy (anonymous):

no.. it's the height of the tetrahedron

OpenStudy (anonymous):

Ok got it.. :) Then your way is correct. :D

OpenStudy (anonymous):

yep :D

OpenStudy (anonymous):

what is the figure and where did 2/3 came from?